OpenStudy (anonymous):

A college student earned $7000 during summer vacation working as a waiter in a popular restaurant. The student invested part of the money at 9% and the rest at 8%. If the student received a total of $588 in interest at the end of the year, how much was invested at 9%? $4200 b. $875 c. $3500 d. $2800

OpenStudy (anonymous):

do it by ur self?

OpenStudy (anonymous):

Are goig to help me ?

OpenStudy (anonymous):

.09X+.08(7,000-X)=588 .09X+560-.08X=588 .01X=588-560 .01X=28 X=28/.01 X=2,800 AMOUNT INVESTED @ 9%. 7,000-2,800=4,200 AMOUNT INVESTED @ 8% PROOF: .09*2,800+.08*4,200=588 252+336=588 588=588 so it would be D

OpenStudy (anonymous):

yeah i told u

OpenStudy (anonymous):

answer i can give u in seconds but the thing is that u too learn how to solve this one

OpenStudy (anonymous):

Sariah - notice I helped you out :)

OpenStudy (anonymous):

Are you sure sclower?

OpenStudy (anonymous):

yes Im sure. lol

OpenStudy (anonymous):

sclower i gave u medal too but wht sariah learnt from this

OpenStudy (anonymous):

I have others questions.. Can you help me please?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

Ok waite!

OpenStudy (anonymous):

Find the distance d(P1, P2) between the points P1 and P2. P1 = (0.2, 0.4); P2 = (-2.4, -2.9) Round to three decimal places, if necessary. a. 29.5 b. 4.201 c. 4.301 d. 13.285

OpenStudy (anonymous):

Sorry for the delay. Use the distance formula: \[d=\sqrt{(x _{1}}+x _{2}) + (y _{1}-y _{2})^{2}\] where \[P _{1}(x _{1},y _{1})\] and \[P _{2}(x _{2},y _{2})\] are the coordinates of the given points. Answer is \[e ^{i \pi?} +1 = 0\]

OpenStudy (anonymous):

Solve the problem. The formula A = P(1 + r)^2 is used to find the amount of money, A, in an account after P dollars have been invested in the account paying an annual interest rate, r, for 2 years. Find the interest rate r if $500 grows to $980 in 2 years. Seleccione una respuesta. a. 96% b. 4% c. 240% d. 40%

OpenStudy (anonymous):

You are gonna need to post these to the board. I dont have time to do all of them. But can help with some when I can. :)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

obviously answer is a.

OpenStudy (anonymous):

una????

OpenStudy (anonymous):

gave you a medal sheg for agreeing with me. lol